OpenStudy (starryknight):

What is the solution to the following system of equations? -8x + 8y = -16 2x + 6y = -28 please help me( please break it down in steps please) free metal for people who help me

OpenStudy (starryknight):

someone please help i really need help on this i dont undrstand this type of equation

OpenStudy (***[isuru]***):

lets take 2x + 6y = -28 .. what will u get if u divide the whole equation by 2 ? i mean... \[\frac{ 2x }{ 2 } + \frac{ 6y }{ 2 } = \frac{ -28 }{ 2 }\]

OpenStudy (starryknight):

-13

OpenStudy (starryknight):

oops i mean -14

OpenStudy (***[isuru]***):

nope... u will get x + 3y = -14 got it ?

OpenStudy (starryknight):

yup

OpenStudy (***[isuru]***):

and lets name x + 3y = -14 ------(1) and lets move to the -8x + 8y = -16 what will u get if u divide the whole expression by 8 ? i mean... \[\frac{ -8x }{ 8 } + \frac{ 8y }{ 8 } = \frac{ -16 }{ 8 }\]

OpenStudy (starryknight):

x+1y= -2 ?

OpenStudy (***[isuru]***):

its should be -x + y = -2 :)

OpenStudy (***[isuru]***):

so... now we have 2 equations first one x + 3y = -14 ----(1) second one -x + y = -2 ----(2) lets add (1) and (2) x + 3y -x + y = -14 -2 what will u get if u simplify this ?

OpenStudy (***[isuru]***):

x + 3y -x + y = -14 -2 if u move like terms close together u can rewrite it as x - x + 3y + y = -16 x and -x will cancel out the x term... which will give u 3y + y = -16 4y = -16 y = -4 got it ?

OpenStudy (starryknight):

i got something but i dont think i got it right

OpenStudy (starryknight):

x+ 3y^2= -14 -2

OpenStudy (starryknight):

oops i mean x^2 + 3y^2=-14-2

OpenStudy (starryknight):

im sorry if i suck i dont know how to do this type of equation

OpenStudy (starryknight):

hello?

OpenStudy (***[isuru]***):

hmmm.... u wont get powers of x or y when u simplify x + 3y -x + y = -14 -2 when u come across something like this... u can add or sub the value in front of like terms... for example in a expression like this 2x + 5x - y 2x and 5x are like terms.. since it contain "x" but 2x and y cant be considered as like terms... since the 2 terms contain "y" and"x" which are 2 separate variables... and when u get 2x + 5x the values in front of each like term is 2 and 5... since the mathematical function u have to do is adding.. u can add 2 to 5 and write 2x + 5x = 7x in the same way... when u come to x + 3y -x + y = -14 -2 x and -x are a pair of like terms while 3y and y is another set of like terms... so if u write the like terms close together... x - x + 3y + y = -14 -2 here if u consider like terms with "x" the mathematical function u have to do is subtract... and the value in front of 2 like terms r 1 and 1... so 1- 1 = 0 therefor x- x = 0 so... u will remain with 3y + y = -14 -2 here the thing u have to do is addition and the values in front of like terms r 3 and 1 so.. 3 + 1 = 4 which means 3y + y = 4y.... so.. u will end up with 4y = -14 -2 = -16 now 4y = -16 if u divide both sides by 4 4y/4 = -16/4 y = -4 got it :) ?

OpenStudy (starryknight):

thank you sooooooooo much

OpenStudy (***[isuru]***):

u r welcome :)